The entrance fee in a carnival is 50 for an adult and 25 for a child. Jun pays 225 for the entrance fee of his family. If his family has equal number of adults and children, how many are they in the family? ( answer is 6 - show your solution )
Answer by srinivas.g(540)

let x be the no of adults int the family
let y be the no of children in the family
family has equal number of adults and children
so , x=y
fee per adult = 50
total fee for x no of adult = 50* x
fee per children = 25
total fee for y no of children = 25* y
Jun pays 225 for the entrance fee of his family
so , 50 x+25y= 225
but x=y
50 x+25 x= 225
75 x =225
divide with 75 on both sides
75 x/75 = 225/75
x =3
no of adults (x)= 3
no of children (y)= 3
total members in the family (x+y)= 3+3 = 6
result is 6 .